### v1

# Additive number theory and the Dyson transform

Abstract.
In 1942 Mann solved a famous problem, the $\alpha+\beta$ conjecture, about the lower bound of the Shnirel’man density of sums of sets of positive integers. In 1945, Dyson generalized Mann’s theorem and obtained a lower bound for the Shnirel’man density of rank $r$ sumsets. His proof introduced a method that evolved into the Dyson transform, an important tool in additive number theory. This paper explains the background of Dyson’s work, gives Dyson’s proof of his theorem, and includes several applications of the Dyson transform, such as Kneser’s inequality for sums of finite subsets of an arbitrary additive abelian group.
Key words and phrases:
Dyson transform, $e$-transform, Mann’s theorem, Shnirel’man density, asymptotic density, $\alpha+\beta$ conjecture, Goldbach conjecture, Kneser’s theorem, Kneser’s inequality.
2020 Mathematics Subject Classification:
11B05,11B13,11B75, 11B77,01A60
MBN supported in part by PSC-CUNY Research Award Program grant 66197-00 54.
1. A theorem on the densities of sets of integers
In 1945 Freeman Dyson [2] published a paper in additive number theory that gave a lower bound for the Shnirel’man density of $r$-rank sums of sets of nonnegative integers. In a commentary on this paper in his Selected Works [3, pp. 5–6], Dyson gave an autobiographical history of this result.
This paper was inspired by …the little book, “Über einige neuere Fortschritte der additiven Zahlentheorie,” by Edmund Landau, [Cambridge University Press, 1937]. As soon as I arrived in Cambridge in 1941, I scoured the book-shops for bargains and bought Landau’s book for six shillings and sixpence…. His Chapter 4 describes the “alpha-beta conjecture”, which was in 1937 a famous unsolved problem. The conjecture concerns the densities of infinite sets of positive integers…. I laid the problem in my heart and worked hard to find a proof….
While I was at Cambridge, I learned that Henry Mann …had proved the alpha-beta conjecture, [Annals of Mathematics, 43, 523–527 (1942)]…. I was sorry that Mann had beaten me in the race to find a proof, but I was glad to see that his strengthening of the conjecture still left me something to do…. The generalization of Mann’s theorem to the sum of more than two sets remained to be proved…. I took the problem with me when I left Cambridge in 1943 to work in the Operational Research Section of the Royal Air Force Bomber Command at High Wycombe. I solved it in the early months of 1944 during a particularly grim period in the history of the Bomber Command…. The Mann Theorem helped to keep me sane amid the insanities of the bombing campaign.
Dyson’s proof of the generalization of Mann’s theorem was inductive, and involved the replacement of one pair $(A_{\ell},A_{n})$ of sets of positive integers with another pair $(A_{\ell}^{\prime},A_{n}^{\prime})$ of sets such that $|A_{n}^{\prime}\cap\{1,\ldots,g\}|<|A_{n}\cap\{1,\ldots,g\}|$ for some positive integer $g$. Through the work of Martin Kneser, a variant of this transformation, variously called the Dyson transform, the $e$-transform, and the Dyson $e$-transform, has become a fundamental tool in additive number theory. This paper explains the mathematical context of Dyson’s paper, presents a proof of Dyson’s addition theorem, and describes some applications of the Dyson transform. For a history of the $\alpha+\beta$ conjecture, see Erdős and Niven [4].
2. Shnirel’man density and sums of primes
Associated to every set $A$ of integers is the number $\sigma(A)$, called the Shnirel’man density of $A$. The Dyson transform was introduced to obtain a lower bound for Shnirel’man density. To appreciate Dyson’s work, we need to understand the importance of Shnirel’man density in additive number theory.
Let $A_{1},A_{2},\ldots,A_{h}$ be nonempty sets of positive integers. The sumset $A_{1}+A_{2}+\cdots+A_{h}$ is the set of all integers that can be written in the form $a_{1}+a_{2}+\cdots+a_{h}$, where $a_{i}\in A_{i}\cup\{0\}$ for all $i=1,2,\ldots,h$. The $h$-fold sumset of a set $A$ is the set
$hA=\underbrace{A+A+\cdots+A}_{\text{$h$ summands}}.$
Let $\mathbf{N}_{0}=\{0,1,2,\ldots\}$ be the set of nonnegative integers. The set $A$ of positive integers is a basis of order $h$ if every nonnegative integer is a sum of $h$ elements of $A\cup\{0\}$, or, equivalently, if $hA=\mathbf{N}_{0}$. The set $A$ is a basis if $A$ is a basis of order $h$ for some positive integer $h$. For example, by Lagrange’s theorem, the set of squares is a basis of order 4. Waring’s problem states that, for every integer $k\geq 2$, there is a positive integer $g(k)$ such that the set of positive $k$th powers is a basis of order $g(k)$.
Let $\mathbf{P}=\{2,3,5,7,\ldots\}$ be the set of prime numbers and let $\widehat{\mathbf{P}}=\{1\}\cup\mathbf{P}$. The Goldbach conjecture that every even integer greater than 2 is the sum of two primes implies that the set $\widehat{\mathbf{P}}$ is a basis of order 3. In 1930, in an extraordinarily original and powerful work, the Soviet mathematician Lev Genrikhovich Shnirel’man [20, 22, 12] proved that there exists an integer $h$ such that every integer greater than 1 is the sum of at most $h$ prime numbers. Equivalently, the set $\widehat{\mathbf{P}}$ is a basis. To obtain this result, Shnirel’man introduced a new density, which we call Shnirel’man density, on sets of integers and proved that every set of positive integers with positive Shnirel’man density is a basis. He used the Brun sieve to prove that the sumset $h^{\prime}\widehat{\mathbf{P}}$ has positive Shnirel’man density for some positive integer $h^{\prime}$, and, mirabile dictu, obtained his result.
Recent results on the Goldbach conjecture include Ramaré’s proof [19] that every even integer $n\geq 2$ is the sum of at most 6 primes, Tao’s proof [23] that every odd integer $n\geq 3$ is the sum of at most 5 primes, and Helfgott’s proof [6] that every odd integer $n\geq 7$ is the sum of 3 primes.
To define Shnirel’man density, we use the counting function $A(x)$ of a set $A$ of integers. This function counts the number of positive integers in $A$ that do not exceed $x$, that is,
$A(x)=\sum_{\begin{subarray}{c}a\in A\\ 1\leq a\leq x\end{subarray}}1.$
The Shnirel’man density of the set $A$ is
$\sigma(A)=\inf_{n\in\mathbf{N}}\frac{A(n)}{n}.$
Shnirel’man density has some usual and some unusual properties. We have $0\leq\sigma(A)\leq 1$ for all sets $A$. The set of odd integers has Shnirel’man density $1/2$. However, $\sigma(A)=0$ if $1\notin A$ and so the set of even integers has Shnirel’man density 00. A fundamental property of Shnirel’man density is that $\sigma(A)=1$ if and only if $A$ contains all positive integers. Thus, the set $A$ is basis of order $h$ for some positive integer $h$ if and only if $\sigma(hA)=1$.
Let $A$ be a set of positive integers. For additive number theory, the key property of Shnirel’man density is that $A$ is a basis of order $h$ for some $h$ if $A$ has positive Shnirel’man density. The proof begins with a simple counting argument.
If $\sigma(A)=\alpha\geq 1/2$, then, for every positive integer $n$, we have
$A(n)\geq\sigma(A)=\alpha n\geq\frac{n}{2}.$
Let
$1\leq a_{1}<\cdots<a_{k}\leq n$
be the $k=A(n)$ positive integers in $A$ that do not exceed $n$. If $n\in A$, then $n=0+n\in 2A$. If $n\notin A$, then $a_{k}\leq n-1$ and so
$11\leq a_{1}<\cdots<a_{k}\leq n-$ 1
and
$11\leq n-a_{k}<\cdots<n-a_{1}\leq n-$ 1 .
Thus, the set $\{1,\ldots,n-1\}$ contains the subsets $A_{1}=\{a_{i}:i=1,\ldots,k\}$ and $A_{2}=\{n-a_{j}:j=1,\ldots,k\}$. Both sets have cardinality $k=A(n)\geq n/2$, and so they cannot be disjoint. Thus, there exist $i,j\in\{1,\ldots,k\}$ such that
$a_{i}=n-a_{j}\in A_{1}\cap A_{2}$
and $n=a_{i}+a_{j}\in 2A$. Thus, $\sigma(A)\geq 1/2$ implies $\sigma(2A)=1$ and so $A$ is a basis of order 2.
The crux of Shnirel’man’s proof is the following lower bound for the density of a sumset. For all sets $A$ and $B$ of integers with $0\in A\cap B$, if $\sigma(A)=\alpha$ and $\sigma(B)=\beta$, then
(1) $\sigma(A+B)\geq\alpha+\beta-\alpha\beta.$
(For a proof of this inequality, see Nathanson [16, Section 7.4].) Equivalently,
$1-\sigma(A+B)\leq(1-\alpha)(1-\beta).$
It follows by induction that if $A_{1},\ldots,A_{h}$ are sets of integers with $\sigma(A_{i})=\alpha_{i}$ for all $i=1,\ldots,h$, then
$1-\sigma\left(A_{1}+\cdots+A_{h}\right)\leq\prod_{i=1}^{h}(1-\alpha_{i}).$
Let $A$ be a set of integers with $\sigma(A)=\alpha$. Applying (1) with $A_{i}=A$ and $\alpha_{i}=\alpha$ for all $i=1,\ldots,h$, we obtain
$1-\sigma\left(hA\right)\leq(1-\alpha)^{h}.$
Equivalently,
$\sigma\left(hA\right)\geq 1-(1-\alpha)^{h}.$
If $\alpha>0$, then $\lim_{h\rightarrow\infty}(1-\alpha)^{h}=0$ and so $\sigma(hA)\geq 1/2$ for all $h\geq h_{0}$ and so $\sigma(2hA)=1$ for all $h\geq 2h_{0}$. Thus, to prove that a set $A$ is a basis, it suffices to prove that the sumset $hA$ has positive Shnirel’man density for some positive integer $h$.
Let $A$ and $B$ be sets of positive integers and let $\sigma(A)=\alpha$ and $\sigma(B)=\beta$. Khinchin [7, 8] conjectured that the Shnirel’man’s density inequality (1) can be strengthened as follows:
(2) $\sigma(A+B)\geq\min(1,\alpha+\beta).$
Let $A+B=C$. Henry B. Mann [14] proved Khinchin’s conjecture in 1942. He wrote, “Stripped of its transcendental content (2) states that
$\frac{C(n)}{n}=1\quad\text{or}\quad\geq\min\frac{A(\ell)}{\ell}+\frac{B(m)}{m},\qquad 1\leq m\leq n,\quad 1\leq\ell\leq n.$
We propose to prove the following sharper theorem.”
Theorem 1 (Mann [14]).
Let $A$ and $B$ be sets of positive integers and let $A+B=C$. Then
$\frac{C(n)}{n}=1\qquad\text{or}\qquad\frac{C(n)}{n}\geq\min_{\begin{subarray}{c}1\leq m\leq n\\ m\notin C\end{subarray}}\frac{A(m)+B(m)}{m}.$
The proof of this sharper result, which Mann called the “Fundamental Theorem,” is elementary but difficult.
Corollary 1.
Let $A$ and $B$ be sets of positive integers and let $A+B=C$. If
$A(m)+B(m)\geq\gamma m$
for all $m=1,\ldots,n$, then
$C(n)\geq\min(1,\gamma)n.$
Proof.
Theorem 1 implies that either $C(n)/n=1$ or $C(n)/n\geq\gamma$, and so $C(n)\geq\min(1,\gamma)n$. ∎
We immediately obtain the $\alpha+\beta$ conjecture.
Corollary 2.
Let $A$ and $B$ be sets of positive integers and let $C=A+B$. If $\sigma(A)=\alpha$ and $\sigma(B)=\beta$, then
$\sigma(A+B)\geq\min(1,\alpha+\beta).$
Proof.
If $\sigma(A)=\alpha$ and $\sigma(B)=\beta$, then, for all $m=1,2,\ldots$, we have $A(m)\geq\alpha m$ and $B(M)\geq\beta m$ and so $A(m)+B(m)\geq(\alpha+\beta)m$. ∎
In 1943, Emil Artin and Peter Scherk [1] improved Mann’s theorem. They wrote, “We found that his method can be simplified considerably and even yields some stronger results.” In 1945, Freeman Dyson [2] refined the Artin-Scherk method and extended Mann’s $\alpha+\beta$ theorem to $r$-fold sums of $n$ sets of positive integers. Like Mann’s original proof, Dyson’s proof is elementary but difficult. An important part of Dyson’s paper was the introduction of what, through the work of Martin Kneser [9], became the Dyson transform.
3. Dyson’s addition theorem
We begin with Dyson’s proof of his addition theorem for the Shnirel’man density of sumsets.
Let $A_{1},\ldots,A_{n}$ be a finite sequence of sets of positive integers. These sets are not necessarily distinct. Let $\mathcal{I}(r)$ be the set of all subsets $I$ of $\{1,\ldots,n\}$ of size $r$. A sum of rank $r$ of the sets $A_{1},\ldots,A_{n}$ is a sumset of the form
$S_{I}=A_{i_{1}}+\cdots+A_{i_{r}}$
where $I=\{i_{1},\ldots,i_{r}\}$ is an $r$-element subset of $\{1,\ldots,n\}$, that is, $I\in\mathcal{I}(r)$. There are $\binom{n}{r}$ subsets $I$ of $\{1,\ldots,n\}$ of size $r$, and so there are $\binom{n}{r}$ sums of rank $r$ of the $n$ sets $A_{1},\ldots,A_{n}$. For every positive integer $m$, we define the rank $r$ counting function
$\phi_{r}(m)=\sum_{I\in\mathcal{I}(r)}S_{I}(m).$
Note that
$\phi_{1}(m)=A_{1}(m)+A_{2}(m)+\cdots+A_{n}(m)$
and
$\phi_{n}(m)=(A_{1}+A_{2}+\cdots+A_{n})(m).$
Theorem 2 (Dyson [2]).
Let $A_{1},\ldots,A_{n}$ be sets of positive integers and let
(3) $\phi_{r}(m)=\sum_{I\in\mathcal{I}(r)}S_{I}(m)$
where the sum is over all sums of rank $r$ of the sets $A_{1},\ldots,A_{n}$. If
$\phi_{1}(m)\geq\gamma m$
for all $m=1,\ldots,g$, then
(4) $\phi_{r}(m)\geq\binom{n-1}{r-1}\min(1,\gamma)m$
for all $m=1,\ldots,g$ and $r=1,\ldots,n$.
This is Dyson’s generalization of Mann’s theorem. It reduces to Mann’s theorem for $n=r=2$.
Proof.
The proof is by double induction on $n$ and $A_{n}(g)$. If $n=1$, then $r=1$ and the theorem is true.
Let $n\geq 2$ and assume that the theorem is true for $n-1$ sets. Let $\phi^{(n-1)}_{r}(m)$ be the rank $r$ counting function of the sets $A_{1},\ldots,A_{n-1}$. By the induction hypothesis, if
$\phi^{(n-1)}_{1}(m)\geq\gamma m$
for all $m=1,\ldots,g$, then
$\phi^{(n-1)}_{r}(m)\geq\binom{n-2}{r-1}\min(1,\gamma)m$
for all $m=1,\ldots,g$ and $r=1,\ldots,n-1$. If $S$ is a set of rank $r$ of the sets $A_{1},\ldots,A_{n}$, then either $S$ is a set of rank $r$ of the sets $A_{1},\ldots,A_{n-1}$ or $S=S^{\prime}+A_{n}$, where $S^{\prime}$ is a set of rank $r-1$ of the sets $A_{1},\ldots,A_{n-1}$. If $g$ is a positive integer such that $A_{n}(g)=0$, then $\phi_{1}(m)=\phi^{(n-1)}_{1}(m)\geq\gamma m$ and $S(m)=S^{\prime}(m)$ for all $m=1,\ldots,g$ and so
$\phi_{r}(m)$ $=\phi^{(n-1)}_{r}(m)+\phi^{(n-1)}_{r-1}(m)$
$\geq\left(\binom{n-2}{r-1}+\binom{n-2}{r-2}\right)\min(1,\gamma)m$
$=\binom{n-1}{r-1}\min(1,\gamma)m$
for all $m=1,\ldots,g$.
Let $g$ be a positive integer such that
$A_{n}(g)\geq 1.$
A Dyson triple is a triple $(a,\ell,c)$ of integers such that
  1. (i)
    $c\in A_{n}\operatorname{\qquad\text{and}\qquad}1\leq c\leq g$
  2. (ii)
    $11\leq\ell\leq n-$ 1
  3. (iii)
    $a\in A_{\ell}\cup\{0\}\qquad\text{or}\qquad a>g$
  4. (iv)
    $a+c\notin A_{\ell}\qquad\text{or}\qquad a+c>g.$
Dyson triples exist. For example, let $c$ be any element of $A_{n}$ such that $1\leq c\leq g$. Such integers $c$ exist because $A_{n}(g)\geq 1$. Choose any positive integer $\ell\leq n-1$. This is possible because $n\geq 2$. Choose any integer $a>g$. Then $a+c>g$ and $(a,\ell,c)$ is a Dyson triple.
Let $a_{0}$ be the smallest integer such that $(a_{0},\ell,c)$ is a Dyson triple for some integers $\ell$ and $c$. Fix an integer $\ell$ such that there is a Dyson triple of the form $(a_{0},\ell,c)$. Consider the set
$T$ =$=\{c\in A_{n}:(a_{0},\ell,c)\text{ is a Dyson triple}\}$
(5) =$=\{c\in A_{n}\cap\{1,\ldots,g\}:c+a_{0}\notin A_{\ell}\text{ or }c+a_{0}>g\}.$
Because there is at least one Dyson triple of the form $(a_{0},\ell,c)$, the set $T$ is a nonempty subset of $\{1,\ldots,g\}$. Equivalently, $T(g)\geq 1$.
Let
(6) $A^{\prime}_{n}$ $=A_{n}\setminus T\hskip 28.45274pt=\{c\in A_{n}:c+a_{0}\in A_{\ell}\text{ and }c+a_{0}\leq g\}$
(7) $A^{\prime}_{\ell}$ $=A_{\ell}\cup(T+a_{0})=A_{\ell}\cup\left\{c+a_{0}:c\in T\right\}$
(8) $A^{\prime}_{i}$ $=A_{i}\qquad\hskip 42.67912pt\text{for all $i\neq\ell,n$ }.$
Because $T$ is a subset of $A_{n}$ with $T(g)\geq 1$, we have
$A^{\prime}_{n}(g)<A_{n}(g)$
and so, by the induction hypothesis, Theorem 2 is true for the $n$ sets $A^{\prime}_{1},\ldots,A^{\prime}_{n}$. Thus, if
$\phi^{\prime}_{1}(m)\geq\gamma m$
for all $m=1,\ldots,g$, then
(9) $\phi^{\prime}_{r}(m)\geq\binom{n-1}{r-1}\min(1,\gamma)m$
for all $m=1,\ldots,g$.
The proof of Dyson’s theorem is an immediate consequence of the following three lemmas.
Lemma 1.
Let $S$ be a set of positive integers and let
(10) $1\leq h\leq g.$
If
(11) $h\in S+A^{\prime}_{\ell}\operatorname{\qquad\text{and}\qquad}h\notin S+A_{\ell}$
then
(12) $h-a_{0}\in S+A_{n}\operatorname{\qquad\text{and}\qquad}h-a_{0}\notin S+A^{\prime}_{n}.$
If
(13) $h\in S+A^{\prime}_{\ell}+A^{\prime}_{n}$
then
(14) $h\in S+A_{\ell}+A_{n}.$
Proof.
We have $A^{\prime}_{\ell}\setminus A_{\ell}\subseteq T+a_{0}$. If $h\in\left(S+A^{\prime}_{\ell}\right)\setminus\left(S+A_{\ell}\right)$, then
$h=s+c+a_{0}$
for some $s\in S\cup\{0\}$ and $c\in T$. Moreover,
$1\leq a_{0}\leq h\leq g.$
It follows that
$h-a_{0}=s+c\in S+T\subseteq S+A_{n}.$
If $h-a_{0}\in S+A^{\prime}_{n}$, then
$h-a_{0}=s+a_{n}$
for some $s\in S\cup\{0\}$ and $a_{n}\in A^{\prime}_{n}\cup\{0\}$.
If $a_{n}\in A^{\prime}_{n}=A_{n}\setminus T$, then
$a_{n}+a_{0}\leq h\leq g\operatorname{\qquad\text{and}\qquad}a_{n}+a_{0}\in A_{\ell}.$
This implies
$h=s+a_{n}+a_{0}\in S+A_{\ell}$
which is absurd.
If $a_{n}=0$, then $a_{n}+a_{0}=a_{0}\in A_{\ell}$ and
$h=s+a_{n}+a_{0}\in S+A_{\ell}$
which is also absurd. Therefore, $h-a_{0}\notin S+A^{\prime}_{n}$. Thus, (11) implies (12).
Suppose that
$h\in A^{\prime}_{n}+A^{\prime}_{\ell}\operatorname{\qquad\text{and}\qquad}h\notin A_{n}+A_{\ell}.$
The relation $A^{\prime}_{n}\subseteq A_{n}$ implies
$h\in A_{n}+A^{\prime}_{\ell}\operatorname{\qquad\text{and}\qquad}h\notin A^{\prime}_{n}+A_{\ell}.$
Applying (11) and (12) with $S=A_{n}$, we obtain
$h-a_{0}\notin A_{n}+A^{\prime}_{n}.$
Applying (11) and (12) with $S=A^{\prime}_{n}$, we obtain
$h-a_{0}\in A_{n}+A^{\prime}_{n}.$
This is absurd, and so
$h\in A^{\prime}_{n}+A^{\prime}_{\ell}\qquad\text{implies}\qquad h\in A_{n}+A_{\ell}.$
It follows that
$h\in S+A^{\prime}_{n}+A^{\prime}_{\ell}\qquad\text{implies}\qquad h\in S+A_{n}+A_{\ell}.$
and so (13) implies (14). This completes the proof. ∎
Lemma 2.
Let
$\phi_{r}(m)=\sum_{I\in\mathcal{I}(r)}S_{I}(m)$
be the rank $r$ counting function for the sets $A_{1},\ldots,A_{n}$ and let
$\phi^{\prime}_{r}(m)=\sum_{I\in\mathcal{I}(r)}S^{\prime}_{I}(m)$
be the rank $r$ counting function for the sets $A^{\prime}_{1},\ldots,A^{\prime}_{n}$. Then
$\phi^{\prime}_{r}(m)\leq\phi_{r}(m)$
for all $m=1,\ldots,g$ and $r=1,\ldots,n$.
Proof.
Let $\ell\in\{1,2,\ldots,n-1\}$ and let $\mathcal{I}^{\prime}(r)$ be the set of all subsets of size $r$ of the set
$\{1,\ldots,\ell-1,\ell+1,\ldots,n-1\}.$
For all positive integers $m$, we have
(15) $\phi_{r}(m)=$ $\sum_{I\in\mathcal{I}^{\prime}(r)}S_{I}(m)+\sum_{I\in\mathcal{I}^{\prime}(r-1)}\left((S_{I}+A_{\ell})(m)+(S_{I}+A_{n})(m)\right)$
$+\sum_{I\in\mathcal{I}^{\prime}(r-2)}(S_{I}+A_{\ell}+A_{n})(m).$
Similarly,
(16) $\phi^{\prime}_{r}(m)=$ $\sum_{I\in\mathcal{I}^{\prime}(r)}S_{I}(m)+\sum_{I\in\mathcal{I}^{\prime}(r-1)}\left((S_{I}+A^{\prime}_{\ell})(m)+(S_{I}+A^{\prime}_{n})(m)\right)$
$+\sum_{I\in\mathcal{I}^{\prime}(r-2)}(S_{I}+A^{\prime}_{\ell}+A^{\prime}_{n})(m).$
It follows from Lemma 1 that
(17) $(S_{I}+A^{\prime}_{\ell})(m)+(S_{I}+A^{\prime}_{n})(m)\leq(S_{I}+A_{\ell})(m)+(S_{I}+A_{n})(m)$
for all $I\in\mathcal{I}^{\prime}(r-1)$ and $m=1,\ldots,g$ and
(18) $\left(S_{I}+A^{\prime}_{\ell}+A^{\prime}_{n}\right)(m)\leq\left(S_{I}+A_{\ell}+A_{n}\right)(m)$
for all $I\in\mathcal{I}^{\prime}(r-2)$ and $m=1,\ldots,g$. Equations (15) and (16) and inequalities (17) and (18) imply
(19) $\phi_{r}^{\prime}(m)\leq\phi_{r}(m)$
for all $m=1,\ldots,g$ and $r=1,\ldots,n$. This completes the proof of Lemma 2. ∎
Lemma 3.
Consider the rank 1 counting functions
$\phi_{1}(m)=\sum_{i=1}^{n}A_{i}(m)$
and
$\phi^{\prime}_{1}(m)=\sum_{i=1}^{n}A^{\prime}_{i}(m).$
If
$\phi_{1}(m)\geq\gamma m$
for all $m=1,\ldots,g$, then
$\phi^{\prime}_{1}(m)\geq\min(1,\gamma)m$
for all $m=1,\ldots,g$
Proof.
For fixed $\ell\in\{1,\ldots,n-1\}$ and $a_{0}\in A_{\ell}$, we have the sets
$T=\{c\in A_{n}:1\leq c\leq g\text{ and }c+a_{0}\notin A_{\ell}\text{ or }c+a_{0}>g\}$
and
$A^{\prime}_{n}$ $=A_{n}\setminus T$
$A^{\prime}_{\ell}$ $=A_{\ell}\cup(T+a_{0})$
$A^{\prime}_{i}$ $=A_{i}\qquad\text{ for all $i\neq\ell,n$.}$
For all $m=1,\ldots,g$, we have, because $T\subseteq A_{n}$,
$A^{\prime}_{n}(m)=A_{n}(m)-T(m)$
and also
$A^{\prime}_{i}(m)=A_{i}(m)$
for all $i\neq\ell,n$.
If $c\in T$ and $c\leq m-a_{0}$, then $c+a_{0}\in T+a_{0}\subseteq A^{\prime}_{\ell}$. Moreover, $c+a_{0}\leq m\leq g$ implies $c+a_{0}\notin A_{\ell}$. Thus, $c+a_{0}\in A^{\prime}_{\ell}\setminus A_{\ell}$. Conversely, if $x\in A^{\prime}_{\ell}\setminus A_{\ell}$ and $x\leq m\leq g$, then $x\in T+a_{0}$ and so $x=c+a_{0}$ for some $c\in T$ with $c\leq m-a_{0}$. It follows that
$A^{\prime}_{\ell}(m)=A_{\ell}(m)+T(m-a_{0}).$
Because
$T(m)\geq T(m-a_{0})$
we have
$\phi^{\prime}_{1}(m)$ $=\sum_{i=1}^{n}A^{\prime}_{i}(m)=\sum_{i=1}^{n}A_{i}(m)-T(m)+T(m-a_{0})$
$=\phi_{1}(m)-T(m)+T(m-a_{0})\leq\phi_{1}(m).$
If $T(m)=T(m-a_{0})$, then
(20) $\phi^{\prime}_{1}(m)=\phi_{1}(m)\geq\gamma m\geq\min(1,\gamma)m.$
Suppose that
$T(m)>T(m-a_{0}).$
In this case we have $\phi^{\prime}_{1}(m)<\phi_{1}(m)$. Moreover, because $T$ is a subset of $A_{n}$, we have
$1\leq T(m)-T(m-a_{0})\leq A_{n}(m)-A_{n}(m-a_{0}).$
Let $b$ be the smallest integer in $A_{n}$ such that
(21) $m-a_{0}<b\leq m\leq g.$
Then $A_{n}(b-1)=A_{n}(m-a_{0})$ and
$\phi^{\prime}_{1}(m)$ $=\phi_{1}(m)-T(m)+T(m-a_{0})$
$\geq\phi_{1}(m)-A_{n}(m)+A_{n}(m-a_{0})$
$=\phi_{1}(m)-A_{n}(m)+A_{n}(b-1).$
We now apply (for the first and only time) the minimality condition that determined the integers $\ell\in\{1,\ldots,n-1\}$ and $a_{0}\in A_{\ell}$ that were used to construct the set $T$. Let $z$ and $c$ be integers such that
$c\in A_{n}$
and
(22) $z-a_{0}<c\leq z\leq g.$
Let $i\in\{1,\ldots,n-1\}$. Let $a$ be an integer in $A_{i}\cup\{0\}$ such that
$a\leq z-c<a_{0}.$
Then
$a+c\leq z\leq g.$
If $a+c\notin A_{i}$, then $(a,i,c)$ is a Dyson triple with $a<a_{0}$, which contradicts the minimality of $a_{0}$. Therefore, $a+c\in A_{i}$ and $c\leq a+c\leq z$ for all $a\in A_{i}\cup\{0\}$ with $a\leq z-c$. There are $A_{i}(z-c)+1$ integers $a\in A_{i}\cup\{0\}$ with $a\leq z-c$. It follows that
(23) $A_{i}(z)\geq A_{i}(z-c)+1+A_{i}(c-1)$
for all $c\in A_{n}$ that satisfy inequality (22).
Inequality (21) allows us to choose $z=m$ and $c=b$ in (22) and (23). We obtain
$A_{i}(m)\geq A_{i}(m-b)+1+A_{i}(b-1)$
for all $i\in\{1,\ldots,n-1\}$. Therefore,
$\phi^{\prime}_{1}(m)$ $\geq\phi_{1}(m)-A_{n}(m)+A_{n}(b-1)$
$=\sum_{i=1}^{n-1}A_{i}(m)+A_{n}(b-1)$
$\geq\sum_{i=1}^{n-1}\left(A_{i}(m-b)+1+A_{i}(b-1)\right)+A_{n}(b-1)$
$=\phi_{1}(b-1)+\phi_{1}(m-b)-A_{n}(m-b)+n-1.$
If $\phi_{1}(m)\geq\gamma m$ for all $m=1,\ldots,g$, then
(24) $\phi^{\prime}_{1}(m)\geq\gamma(b-1)+\phi_{1}(m-b)-A_{n}(m-b)+n-1.$
Let
$\delta=\min(1,\gamma).$
We shall prove that there exists no positive integer $z$ such that
(25) $z<a_{0},\qquad z\leq g,\operatorname{\qquad\text{and}\qquad}\phi_{1}(z)-A_{n}(z)<\delta z.$
If such integers exist, let $z$ be the smallest one. We have
(26) $A_{n}(z)>\phi_{1}(z)-\delta z\geq\gamma z-\delta z\geq 0.$
Because $A_{n}(z)>0$, there exists $c\in A_{n}$ with $c\leq z$. We have
$z-a_{0}<0<c\leq z\leq g$
and so $c$ and $z$ satisfy inequality (22). Applying inequality (23), we obtain
$\phi_{1}(z)-A_{n}(z)$ $=\sum_{i=1}^{n-1}A_{i}(z)$
$\geq\sum_{i=1}^{n-1}\left(A_{i}(z-c)+1+A_{i}(c-1)\right)$
(27) $=\phi_{1}(z-c)-A_{n}(z-c)+\phi_{1}(c-1)-A_{n}(c-1)+n-1.$
Both $z-c$ and $c-1$ are less than $z$, which is the smallest number satisfying (25), and so
$\phi_{1}(z-c)-A_{n}(z-c)\geq\delta(z-c)$
and
$\phi_{1}(c-1)-A_{n}(c-1)\geq\delta(c-1).$
Inserting these inequalities into (3) gives
(28) $\phi_{1}(z)-A_{n}(z)\geq\delta(z-1)+n-1.$
Because $n\geq 2$ and $\delta\leq 1$, inequalities (26) and (28) imply
$\delta z>\phi_{1}(z)-A_{n}(z)\geq\delta z+n-1-\delta\geq\delta z$
which is absurd. Therefore, no integer $z$ satisfies inequalities (25).
Recall that $b$ is the smallest integer in the set $A_{n}$ such that
$m-a_{0}<b\leq m\leq g$
and so
$m-b<a_{0}\operatorname{\qquad\text{and}\qquad}m-b\leq g.$
The impossibility of (25) implies
$\phi_{1}(m-b)-A_{n}(m-b)\geq\delta(m-b).$
Because $\delta=\min(1,\gamma)$, inequality (24) gives
$\phi^{\prime}_{1}(m)$ $\geq\gamma(b-1)+\phi_{1}(m-b)-A_{n}(m-b)+n-1$
$\geq\gamma(b-1)+\delta(m-b)+n-1$
$=\delta m+(\gamma-\delta)(b-1)+n-1-\delta$
$\geq\delta m=\min(1,\gamma)m.$
Combining this inequality with (20), we obtain $\phi^{\prime}_{1}(m)\geq\min(1,\gamma)m$ for all $m=1,\ldots,g$. This completes the proof of Lemma 3. ∎
By Lemma 3, if $\varphi_{1}(m)\geq\gamma m$ for $m=1,\ldots,g$, then $\varphi^{\prime}_{1}(m)\geq\min(1,\gamma)m$ for $m=1,\ldots,g$. Applying the induction hypothesis (9) and Lemma 2, we obtain
$\phi_{r}(m)\geq\phi^{\prime}_{r}(m)\geq\binom{n-1}{r-1}\delta m$
for all $m=1,\ldots,g$ and $r=1,\ldots,n$. This completes the proof of Dyson’s Theorem. ∎
4. Asymptotic density and Kneser’s theorem
Let $A$ and $B$ be sets of positive integers. By Mann’s $\alpha+\beta$ theorem, Shnirel’man density is superadditive in the following sense:
$\sigma(A+B)\geq\min(1,\sigma(A)+\sigma(B)).$
The lower asymptotic density of the set $A$ is
$d_{L}(A)=\liminf_{n=1,2,\ldots}\frac{A(n)}{n}.$
Asymptotic density is not superadditive. For example, let $k$, $\ell$, and $m$ be positive integers such that
$k+\ell\leq m.$
Consider the sets
$A=\left\{n\in\mathbf{N}_{0}:n\equiv r\pmod{m}\text{ for some }r\in\{0,1,\ldots,k-1\}\right\}$
and
$B=\left\{n\in\mathbf{N}_{0}:n\equiv s\pmod{m}\text{ for some }s\in\{0,1,\ldots,\ell-1\}\right\}.$
We have
$d_{L}(A)=\frac{k}{m}\operatorname{\qquad\text{and}\qquad}d_{L}(B)=\frac{\ell}{m}.$
The sumset
$A+B=\left\{n\in\mathbf{N}_{0}:n\equiv t\pmod{m}\text{ for some }t\in\{0,1,\ldots,k+\ell-2\}\right\}.$
has lower asymptotic density
$d_{L}(A+B)$ $=\frac{k+\ell-1}{m}$
$=d_{L}(A)+d_{L}(B)-\frac{1}{m}$
$<d_{L}(A)+d_{L}(B).$
In 1953, Martin Kneser [9] proved the beautiful theorem that the only counterexamples to superadditivity for lower asymptotic density of sumsets are pairs of sets $A$ and $B$ that are essentially unions of congruence classes modulo $m$ for some positive integer $m$. The long and difficult proof is a tour de force based on a masterful application of iterated Dyson transforms. For proofs of Kneser’s theorem, see Kneser [9] and also Halberstam-Roth [5], Mann [15], and Nathanson [18].
5. The $e$-transform in an abelian group
Analogous to his theorem on the lower asymptotic density of sums of sets of integers, Kneser also proved a theorem on sums of finite subsets of an abelian group. Here is a simple form of the Dyson transform. Let $G$ be an abelian group, written additively, with identity element 00. For $X\subseteq G$ and $e\in G$, we define the translates
$X+e=\{x+e:x\in X\}$
and
$X-e=\{x-e:x\in X\}.$
Let $(A,B)$ be a pair of nonempty subsets of $G$. For all $e\in A$, let
$A(e)=A\cup(B+e)$
and
$B(e)=B\cap(A-e)$
We have
$A\subseteq A(e)$
and
$B(e)\subseteq B.$
The pair of sets $(A(e),B(e))$ is the $e$-transform of the pair $(A,B)$.
Theorem 3.
Let $A$ and $B$ be nonempty subsets of the abelian group $G$ and let $e\in A$. Then
(29) $A(e)+B(e)\subseteq A+B$
and
(30) $A(e)\setminus A=e+\left(B\setminus B(e)\right).$
If $A$ and $B$ are finite sets, then
(31) $|A|+|B|=|A(e)|+|B(e)|.$
If $e\in A$ and $0\in B$, then $e\in A(e)$ and $0\in B(e)$.
Proof.
Let $a^{\prime}+b^{\prime}\in A(e)+B(e)$, where $a^{\prime}\in A(e)$ and $b^{\prime}\in B(e)$. Then $a^{\prime}\in A$ or $a^{\prime}\in B+e$. If $a^{\prime}\in A$, then $b^{\prime}\in B(e)\subseteq B$ implies $a^{\prime}+b^{\prime}\in A+B$. If $a^{\prime}\in B+e$, then $a^{\prime}=b^{\prime\prime}+e$ for some $b^{\prime\prime}\in B$. We have $b^{\prime}\in B(e)$ and so $b^{\prime}=a^{\prime\prime}-e$ for some $a^{\prime\prime}\in A$. It follows that
$a^{\prime}+b^{\prime}=(b^{\prime\prime}+e)+(a^{\prime\prime}-e)=a^{\prime\prime}+b^{\prime\prime}\in A+B.$
This proves (29).
We have
$b\in B\setminus B(e)$
if and only if
$b\in B\operatorname{\qquad\text{and}\qquad}b\notin A-e$
if and only if
$b+e\in B+e\operatorname{\qquad\text{and}\qquad}b+e\notin A$
if and only if
$b+e\in A(e)\setminus A.$
This proves (30).
If $A$ and $B$ are finite sets, then (30) implies
$\left|A(e)\right|-\left|A\right|=\left|A(e)\setminus A\right|=\left|B\setminus B(e)\right|=\left|B\right|-\left|B(e)\right|.$
This proves (31). ∎
As an example of the use the $e$-transform in additive number theory, we prove I. Chowla’s generalization of the Cauchy-Davenport theorem. We begin with a simple lemma.
Lemma 4.
Let $G$ be a finite group, written multiplicatively and not necessarily abelian. Let $A$ and $B$ be subsets of $G$ and let
$AB=\{ab:a\in A\text{ and }b\in B\}.$
If
$|A|+|B|>|G|$
then
$AB=G.$
Proof.
Let $x\in G$ and let
$xB^{-1}=\{xb^{-1}:b\in B\}.$
Then
$\left|xB^{-1}\right|=|B|$
and so
$|A|+\left|xB^{-1}\right|=|A|+|B|>|G|.$
By the pigeonhole principle, the sets $A$ and $xB^{-1}$ are not disjoint, and so there exists $a\in A$ and $b\in B$ such that $a=xb^{-1}$. Thus, $x=ab\in AB$ and so $AB=G$. This completes the proof. ∎
Theorem 4 (I. Chowla).
Let $m\geq 2$ and let $A$ and $B$ be nonempty subsets of the additive abelian group $\mathbf{Z}/m\mathbf{Z}$ of congruence classes modulo $m$. If $0\in B$ and $\gcd(b,m)=1$ for all $b\in B\setminus\{0\}$, then
(32) $|A+B|\geq\min(m,|A|+|B|-1).$
Proof.
By Lemma 4, if $|A|+|B|>m$, then $A+B=\mathbf{Z}/m\mathbf{Z}$ and $|A+B|=m$. Thus, we can assume that
$|A|+|B|\leq m.$
Inequality (32) also holds if $|A|=1$ or $|B|=1$, and so we can assume that
$|A|\geq 2\operatorname{\qquad\text{and}\qquad}|B|\geq 2.$
If the theorem is false, then there is a pair of sets $(A,B)$ such that $|A|\geq 2$ and $|B|\geq 2$ and
$|A+B|<|A|+|B|-1.$
Choose the counterexample $(A,B)$ so that the cardinality of $B$ is minimal. Because $|B|\geq 2$, there exists $b^{*}\in B$ with $b^{*}\neq 0$. We have $\gcd(b^{*},m)=1$.
If $a+b^{*}\in A$ for all $a\in A$, then $a+jb^{*}\in A$ for all $j=0,1,2,3,\ldots$. Because the set $A$ is finite, there exist integers $j_{1}<j_{2}$ with $j_{0}=j_{2}-j_{1}$ minimal such that $a+j_{1}b^{*}=a+j_{2}b^{*}$. It follows that
$j_{0}b^{*}=\left(a+j_{1}b^{*}\right)-\left(a+j_{2}b^{*}\right)=0$
The divisibility condition $\gcd(b^{*},m)=1$ implies $j_{0}=m$ and $A=G$, which is absurd. Therefore, there exists $e\in A$ with $e+b^{*}\notin A$.
We apply this element $e$ to the pair $(A,B)$ to obtain the $e$-transform $(A(e),B(e))$. By Theorem 3,
$|A(e)+B(e)|$ $\leq|A+B|<|A|+|B|-1$
$=|A(e)|+|B(e)|-1.$
Because $e\in A$ and $0\in B$, we have $0\in B(e)=B\cap(A-e)$. Moreover, $B(e)\subseteq B$ implies $\gcd(b,m)=1$ for all $b\in B(e)\setminus\{0\}$. Because $e+b^{*}\notin A$, we have $b^{*}\notin A-e$ and so $b^{*}\notin B(e)$. Thus, $B(e)$ is a proper subset of $B$, which contradicts the minimality of $|B|$. This completes the proof. ∎
From I. Chowla’s theorem we immediately obtain the Cauchy-Davenport theorem.
Theorem 5 (Cauchy-Davenport).
Let $p$ be a prime number and let $A$ and $B$ be nonempty subsets of $\mathbf{Z}/p\mathbf{Z}$. Then
(33) $|A+B|\geq\min(p,|A|+|B|-1).$
Proof.
Let $A$ and $B$ be nonempty subsets of $\mathbf{Z}/p\mathbf{Z}$. Let $b\in B$ and let $B^{\prime}=B-b$. Then $|B^{\prime}|=|B|$ and $|A+B^{\prime}|=|A+B-b|=|A+B|$. Also, $0\in B^{\prime}$.
Because $p$ is prime, we have $\gcd(b^{\prime},p)=1$ for all $b^{\prime}\in B^{\prime}$ with $b^{\prime}\neq 0$. Applying Theorem 4 with $m=p$ to the sets $A$ and $B^{\prime}$, we obtain
$|A+B|=|A+B^{\prime}|\geq\min(p,|A|+|B^{\prime}|-1)=\min(p,|A|+|B|-1).$
This completes the proof. ∎
Martin Kneser proved the following beautiful group theoretic inequality.
Theorem 6 (Kneser).
Let $G$ be an abelian group, written additively, and let $A$ and $B$ be finite nonempty subsets of $G$. If $|A|+|B|\leq|G|$, then there exists a proper subgroup $H$ of $G$ such that
$|A+B|\geq|A|+|B|-|H|.$
Let $p$ be a prime number and let $G=\mathbf{Z}/p\mathbf{Z}$. The only proper subgroup of $\mathbf{Z}/p\mathbf{Z}$ is $H=\{0\}$, and so $|A+B|\geq|A|+|B|-1$. Thus, Kneser’s inequality immediately implies the Cauchy-Davenport theorem (Theorem 5).
Proof.
The proof is by induction on $|B|$. If $|B|=1$, then
$|A+B|=|A|\geq|A|+1-|H|$
for every subgroup $H$ of $G$.
Let $k\geq 2$ and assume that Kneser’s inequality is valid for all pairs $(A,B)$ of nonempty finite subsets of $G$ with $|B|<k$.
Let $(A,B)$ be a pair of nonempty finite subsets of $G$ with $|B|=k$. There are two cases.
In the first case, we have
$a+b-b^{\prime}\in A$
for all $a\in A$ and $b,b^{\prime}\in B$. Then
$A+b-b^{\prime}\subseteq A$
for all $b,b^{\prime}\in B$. Because $A$ is finite and $|A+b-b^{\prime}|=|A|$, it follows that
$A+b-b^{\prime}=A$
for all $b,b^{\prime}\in B$. Let $H$ be the subgroup of $G$ generated by the set $\{b-b^{\prime}:b,b^{\prime}\in B\}$. Note that $A\neq G$ because $|A|+|B|\leq|G|$ and that $|B|\leq|H|$. We have
$A+H=A\neq G$
and so $H\neq G$. Thus, $H$ is a proper subgroup of $G$ such that
$|A+B|\geq|A|\geq|A|+|B|-|H|.$
In the second case, there exist $a\in A$ and $b,b^{\prime}\in B$ such that
$a+b-b^{\prime}\not\in A.$
Let
$e=a-b^{\prime}.$
We apply the $e$-transform to the pair $(A,B)$ to obtain the pair $(A(e),B(e))$, where
$A(e)=A\cup(B+e)$
and
$B(e)=B\cap(A-e).$
We have $b+e=a+b-b^{\prime}\notin A$ and so $b\notin A-e$. Thus, $|A(e)|>|A|$ and $|B(e)|<|B|$. Also, $b^{\prime}=a-(a-b^{\prime})=a-e\in A-e$ and so $b^{\prime}\in B\cap(A-e)\neq\emptyset$. Therefore, $B(e)$ is a nonempty finite subset of $G$ with $|B(e)|<|B|$ and so the induction hypothesis applies to the pair $(A(e),B(e))$. Applying Theorem 3, we obtain
$|A+B|\geq|A(e)+B(e)|\geq|A(e)|+|B(e)|-|H|=|A|+|B|-|H|.$
This completes the proof. ∎
Kneser also proved a more precise version of Theorem 6. The stabilizer of a nonempty subset $X$ of an abelian group $G$ is the set
$H(X)=\{g\in G:X+g=X\}.$
The stabilizer of $X$ is a subgroup of $G$.
Theorem 7 (Kneser).
Let $G$ be an abelian group, written additively, let $A$ and $B$ be finite nonempty subsets of $G$, and let $H=H(A+B)$ be the stabilizer of the sumset $A+B$. If
$|A+B|<|A|+|B|$
then
$|A+B|=|A+H|+|B+H|-|H|.$
Proof.
For Kneser’s proofs, see [10, 11]. See also Nathanson [17, pp. 110–116]. ∎